How many degrees on a circle graph would each be given from the following? (Round your answers to the nearest tenth.)

Wear digital watch 35 % o
Wear traditional watch 57 % o
Wear no watch 8 % o
Answer #1

Solution:

Percentage wearing in digital watch = 35%

Degree in circle graph = (35 * 360/100) = 126 degree

Percentage wearing in traditional watch = 57%

Degree in circle graph = (57 * 360/100) = 205.20 degree

Percentage wear no watch = 8%

Degree in circle graph = (8 * 360/100) = 28.80 degree